OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. m =3/5 , P(3, 2) and Q(8,_)

OpenStudy (anonymous):

would it be 5

OpenStudy (ajprincess):

yep:)

OpenStudy (anonymous):

ok thank you what about this one I am not getting Use the slope formula to find the missing coordinate R(0, _), S(-4, 2) m = 5

Miracrown (miracrown):

That’s correct, good job!

OpenStudy (ajprincess):

m=y_2-y_1/x_2-x_1 5=2-y_1/-4 5*-4=2-y_2 -20=2-y_2 y_2=?

Miracrown (miracrown):